OpenStudy (haleyelizabeth2017):

Solve by sublimation. Can you just tell me what the first few steps are @myininaya? -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

OpenStudy (perl):

sublimation wont help you solve that (sublimation is a chemistry turn, turning solid into gas skipping liquid phase)

OpenStudy (perl):

chemistry term*

myininaya (myininaya):

i think he just combine the words substitution and elimination making a new word (or a new one for me at least) sublimation But maybe it should be sublimination.

OpenStudy (perl):

right, he meant elimination. but i thought it was still funny

myininaya (myininaya):

Anyways... You could try to eliminate the x variable getting two equations in terms of 2 variables y and z. (I thought it was cute and funny too :p)

OpenStudy (anonymous):

To start with: Divide both sides of the third equation by 2. Add the result above to the first equation. y = 6

myininaya (myininaya):

or getting two equations in terms of z and x since the last one is in terms of just z and x but you can try rob's way too

OpenStudy (haleyelizabeth2017):

Sorry! I meant substitution!

OpenStudy (haleyelizabeth2017):

@myininaya I meant substitution :)

myininaya (myininaya):

Ok so first last write your last equation as x+z=2

myininaya (myininaya):

now your first equation is -(x+z)-y=-8

myininaya (myininaya):

right ?

OpenStudy (haleyelizabeth2017):

umm...one second

OpenStudy (haleyelizabeth2017):

I don't remember.....this is just confusing.....I'm used to doing 2 variables, not 3....:(

myininaya (myininaya):

Ok I haven't done anything yet except rewrite your first and last equation

myininaya (myininaya):

First equation is -(x+z)-y=-8 Last equation is x+z=2

OpenStudy (haleyelizabeth2017):

okay :) why do we need to write the first one like that?

myininaya (myininaya):

Do you see how you can sub the last equation into the first?

OpenStudy (haleyelizabeth2017):

-put it in as -(2)-y=8

myininaya (myininaya):

I wrote it like that so it easy for you to tell tat you can replace x+z with 2

myininaya (myininaya):

yep

myininaya (myininaya):

well -2-y=-8*

OpenStudy (haleyelizabeth2017):

okay :) sorry so um y=6

myininaya (myininaya):

yep

OpenStudy (haleyelizabeth2017):

no how do I do the x and z?

OpenStudy (haleyelizabeth2017):

now*

myininaya (myininaya):

So remember we had these equations: \[x+z=2 \\ y=6 \\ -4x+4(6)+5z=7 \] You can solve that first one for x and then replace the x in the last equation with something in terms of z

OpenStudy (haleyelizabeth2017):

with something...? what do you mean? OOHHHH!!!! so I can isolate z or x and plug it into a different equation! I get it now :P

OpenStudy (haleyelizabeth2017):

right?

OpenStudy (haleyelizabeth2017):

@myininaya?

myininaya (myininaya):

yes

myininaya (myininaya):

Well I mean to isolate x in the first equation then take that x and replace in the last equation this will give an equation in terms of z

OpenStudy (haleyelizabeth2017):

okay :) thank you. I'm gonna plug it into the 1st one instead :) It seems easiest to isolate

myininaya (myininaya):

either way

OpenStudy (haleyelizabeth2017):

-4(2-z)+4(6)+5z=7?

myininaya (myininaya):

Beautiful

myininaya (myininaya):

So we have another equation just in terms of z

OpenStudy (haleyelizabeth2017):

z=-1

myininaya (myininaya):

-8+4z+24+5z=7 -8+9z+24=7 9z=15-24 9z=-9 ok yeah I see z=-1

OpenStudy (haleyelizabeth2017):

x=3

myininaya (myininaya):

checking: z+x=2 -1+x=2

myininaya (myininaya):

looks good

myininaya (myininaya):

are you going to check your solution like you did last time?

OpenStudy (haleyelizabeth2017):

yup

myininaya (myininaya):

that's good practice :)

OpenStudy (haleyelizabeth2017):

thanks :)